OpenStudy (anonymous):

If the temperature on the ground is 60 degrees F, then the air temperature x miles high is given T=60-29x. Determine the altitudes at which the air temperature is less than 4.1 degree F. The altitudes at which the air temperature is less then 4.1 degree F are more than ______ miles.

OpenStudy (anonymous):

You are looking for x where T = 60 - 4.1; So 60-4.1 = 60-29x. Just solve for x.

OpenStudy (anonymous):

OK I got as far as 60-4.1+60=60-29x+60 4.1 = 29x

OpenStudy (anonymous):

That is correct, how you just need to divide 4.1 by 29 for the answer in miles.

OpenStudy (anonymous):

Oh gosh...thanks